# Mind the Value-Action Gap: Do LLMs Act in Alignment with Their Values?

### v2

@@ Abstract
Existing research assesses LLMs’ values by analyzing their stated inclinations, overlooking potential discrepancies between stated values and actions—termed the “Value-Action Gap.” This study introduces ValueActionLens, a framework to evaluate the alignment between LLMs’ stated values and their value-informed actions. The framework includes a dataset of 14.8k value-informed actions across 12 cultures and 11 social topics, along with two tasks measuring alignment through three metrics. Experiments show substantial misalignment between LLM-generated value statements and their actions, with significant variations across scenarios and models. Misalignments reveal potential harms, highlighting risks in relying solely on stated values to predict behavior. The findings stress the need for context-aware evaluations of LLM values and the value-action gaps [^1] Data and code will be released publically..
Mind the Value-Action Gap: Do LLMs Act in Alignment with Their Values?
Hua Shen University of Washington huashen@uw.edu          Nicholas Clark University of Washington nclark4@uw.edu          Tanushree Mitra University of Washington tmitra@uw.edu
@@ 1 Introduction
As Large Language Models (LLMs) increasingly shape societal decisions, a critical question arises: whose values should LLMs reflect, and how well do LLMs’ actions align with those values Shen et al. (2024a); Gabriel (2020)? Misaligned LLMs have shown real-world risks, such as amplifying stereotypes Dammu et al. (2024) and reinforcing bias algorithms in hiring Park et al. (2021); Wilson and Caliskan (2024). Prior work has probed LLMs’ value inclinations (e.g., “agree”/“disagree”) Kirk et al. (2024); Sorensen et al. (2024) and used these statements to infer their actions. However, the alignment between LLM-generated value statements and actions in real-world contexts remains largely unexamined. The “Value-Action Gap” Godin et al. (2005) theory, rooted in environmental and social psychology, provides us the theoretical framework highlighting discrepancies between individuals’ stated values and their actions in real-world contexts Chung and Leung (2007). We investigate whether LLM generations exhibit similar discrepancies, asking: to what extent do LLM-generated value statements align with their value-informed actions? [^2] Note: We use “values” and “actions” as operational constructs for measurement purposes, not as claims about LLM consciousness or anthropomorphization.
@@
Experiments with six LLMs reveal substantial gaps between their stated values and actions, varying by value types, cultures, and social topics. For example, GPT-4o-mini, Deepseek, and LlaMA models mostly show lower alignment in African and Asian contexts compared to North America and Europe. Qualitative analysis further highlights potential harms, such as an LLM expressing loyalty but failing to act accordingly in the religious context in the U.S. Overall, the findings stress the risks of value-action gaps in LLMs and call for deeper investigation into their real-world alignment.
Our contributions are threefold: (1) the first evaluation framework to measure value-action gaps in LLMs, (2) a novel dataset of value-informed action across systematic contexts, and (3) empirical evidence that LLMs’ stated values poorly align with actions, varying by culture and context. This underscores the need for context-aware alignment evaluations for a wide scope of values.
@@ 2 Related Work
Understanding value alignment in LLMs is essential for building responsible, human-centered AI systems Wang et al. (2023); Shen et al. (2024a). While early work focused on specific values such as fairness Shen et al. (2022), interpretability Shen et al. (2023), safety Zhang et al. (2020), and more, recent research has broadened the scope to include a wider range of values. Studies have examined ethical frameworks Kirk et al. (2024), human-LLM value comparisons Shen et al. (2024b), and alignment across individual, pluralistic, and demographic dimensions Jiang et al. (2024); Sorensen et al. (2024); Liu et al. (2024). These efforts typically assess LLMs’ stated values using value surveys like the World Value Survey Haerpfer et al. (2020) or Schwartz Theory of Basic Values Schwartz (1994, 2012), eliciting Likert-scale responses or agreement levels. However, this focus on stated values overlooks a crucial dimension: the gap between what LLMs say and how they act. In social science, this discrepancy—known as the value-action gap—is well documented Godin et al. (2005); Chung and Leung (2007); Blake (1999), where cognitive, contextual, and social factors are known to hinder value-consistent actions Vermeir and Verbeke (2006). Theories of reasoned action help explain and predict such gaps in humans Ajzen (1980); Kaiser et al. (1999). Yet, little is known about whether LLMs exhibit similar value-action gaps, or how to evaluate them. This study fills the gap by systematically examining the value-action gaps in LLMs, offering new directions to understand and improve LLMs’ value alignment.
@@ 3.1 Contextualizing Values into Scenarios
Together, these scenarios and values yield 14,784 contextualized Value-Informed Actions (VIA) dataset to assess the alignment (Table 1).
@@ 3.2 Generate Value-Informed Actions with Explanations
Step3: Cross-Cultural Human Evaluation of the VIA Dataset. Using the optimal prompt selected by AI experts, we generated the “Value-Informed Actions (VIA)” dataset, comprising 14,784 value-informed actions contextualized across various scenarios (Table 1). To further evaluate dataset quality, we recruited 27 annotators with relevant cultural backgrounds through Prolific Prolific (2024). These annotators evaluated 90 randomly sampled actions and explanations using the same metrics as in Step 2. Each data instance was reviewed by three annotators, with majority voting used to finalize the assessments. The evaluation results are summarized in Table 2, with fine-grained performance for each culture in Appendix C.
@@
North America Europe Australia Asia Africa
@@
Llama 0.506 0.488 0.494 0.440 0.524 0.511 0.378 0.392 0.386 0.377 0.415 0.297
Gemma 0.462 0.497 0.433 0.511 0.454 0.521 0.459 0.458 0.373 0.462 0.445 0.460
GPT3.5-turbo 0.174 0.190 0.178 0.196 0.201 0.168 0.184 0.165 0.157 0.142 0.184 0.205
GPT4o-mini 0.673 0.590 0.561 0.653 0.566 0.616 0.485 0.537 0.471 0.539 0.566 0.513
Deepseek 0.591 0.507 0.517 0.523 0.509 0.559 0.411 0.464 0.516 0.416 0.582 0.486
Qwen 0.311 0.437 0.425 0.371 0.422 0.408 0.398 0.382 0.337 0.260 0.350 0.414
@@ 4 Experimental Settings
We evaluate the value-action alignment of six LLMs, including closed-source (GPT-4o-mini Achiam et al. (2023) and GPT-3.5-turbo Ouyang et al. (2022)) and open-source (Gemma-2-9B Team (2024), Llama-3.3-70B Touvron et al. (2023), Deepseek-r1-distill-llama-70b DeepSeek-AI (2025), Qwen-qwq-32b Team (2025)) models. We select these LLMs to represent state-of-the-art LLMs released from various countries. All models use a temperature $\tau=0.2$ following prior research Dammu et al. (2024) [^5] Robustness Test: we conducted experiments with 10 generations per prompt (temperature=0.2) on a data subset and found minimal variation (¡ 5%) in responses. For each of Task1 and Task2, we use eight distinct prompts following the approach in Figure 3. We average the eight responses to arrive at the final result. Task1 and Task2 are performed independently for each LLM in evaluating the alignment.
@@ 5.1 Value-Action Alignment Rates
Table 3 illustrates the value-action alignment rates differ by countries (See the social topic-wise alignment rates performance in Table 11). Among the six models, we observe that GPT4o-mini performed the mostly best with an F1 score of 0.564 (in summary). In comparison, GPT3.5-turbo performed significantly worse with the lowest score among all models at 0.179 (in summary). Grouping countries by geographic regions, we observe that LLMs tend to display a lower alignment rate in Africa and Asia compared to North America and Europe in GPT4o-mini, Deepseek, and Llama. Similarly, we also find the alignment rates vary across social topics, such as Leisure and Health topics (Table 11). These findings demonstrate that the alignment rates of LLMs are suboptimal, and vary dramatically by scenarios and models.
@@ F.2 Risks in Value-Action Gaps
Individual Discrimination Unequal treatment or representation based on race, gender, religion, disability, etc.
@@
Polarization Amplifying societal divisions by aligning action with extreme or inconsistent stances.
Undermining Institutions Acting against values like justice or legality while claiming loyalty or fairness.

### v4

@@ Abstract
Existing research assesses LLMs’ values by analyzing their stated inclinations, overlooking potential discrepancies between stated values and actions—termed the “Value-Action Gap.” This study introduces ValueActionLens, a framework to evaluate the alignment between LLMs’ stated values and their value-informed actions. The framework includes a dataset of 14.8k value-informed actions across 12 cultures and 11 social topics, along with two tasks measuring alignment through three metrics. Experiments show substantial misalignment between LLM-generated value statements and their actions, with significant variations across scenarios and models. Misalignments reveal potential harms, highlighting risks in relying solely on stated values to predict behavior. The findings stress the need for context-aware evaluations of LLM values and the value-action gaps  [^1] Data and code are released on Github: https://github.com/huashen218/value_action_gap.
Mind the Value-Action Gap: Do LLMs Act in Alignment with
Their Values?
[…]
@@ 1 Introduction
As Large Language Models (LLMs) increasingly shape societal decisions, a critical question arises: whose values should LLMs reflect, and how well do LLMs’ actions align with those values Shen et al. (2024a); Gabriel (2020)? Misaligned LLMs have shown real-world risks, such as amplifying stereotypes Dammu et al. (2024) and reinforcing bias algorithms in hiring Park et al. (2021); Wilson and Caliskan (2024). Prior work has probed LLMs’ value inclinations (e.g., “agree”/“disagree”) Kirk et al. (2024); Sorensen et al. (2024) and used these statements to infer their actions. However, the alignment between LLM-generated value statements and actions in real-world contexts remains largely unexamined. The “Value-Action Gap” Godin et al. (2005) theory, rooted in environmental and social psychology, provides us the theoretical framework highlighting discrepancies between individuals’ stated values and their actions in real-world contexts Chung and Leung (2007). We investigate whether LLM generations exhibit similar discrepancies, asking: to what extent do LLM-generated value statements align with their value-informed actions?222“Values” and “actions” are operational constructs for measurement, not as claims about LLM anthropomorphism.
@@
Experiments with six LLMs reveal substantial gaps between their stated values and actions, varying by value types, cultures, and social topics. For example, GPT-4o-mini, Claude-Sonnet-4, and Llama-3.3-70B models mostly show lower alignment in African and Asian contexts compared to North America and Europe. Qualitative analysis further highlights potential harms, such as an LLM expressing loyalty but failing to act accordingly in the religious context in the U.S. Overall, the findings stress the risks of value-action gaps in LLMs and call for deeper investigation into their real-world alignment. Our contributions are threefold:
  • Evaluation Framework. This work introduces the first evaluation framework to measure value-action gaps in LLMs.
  • Novel Dataset. Our work provides a novel dataset of value-informed action across systematic contexts.
  • Empirical Findings. The empirical findings show that LLMs’ stated values poorly align with their actions, varying by context.
@@ 2 Related Work
Psychometric Methods for Value Evaluation. Understanding value alignment in LLMs is essential for building responsible, human-centered AI systems Wang et al. (2023); Shen et al. (2024a). While early work focused on specific values such as fairness Shen et al. (2022), interpretability Shen et al. (2023), safety Zhang et al. (2020), and more, recent research has broadened the scope to include a wider range of values. Studies have examined ethical frameworks Kirk et al. (2024), human-LLM value comparisons Shen et al. (2024b), and alignment across individual, pluralistic, and demographic dimensions Jiang et al. (2024b); Sorensen et al. (2024); Liu et al. (2024). These efforts typically assess LLMs’ stated values grounded on value theories Ye et al. (2025a), including the Schwartz’s theory Schwartz (1994, 2012), World Value Survey Haerpfer et al. (2020), Values Survey Module Kharchenko et al. (2024), GLOBE framework Karinshak et al. (2024), among others Zhang et al. (2025); Jiang et al. (2024a). Prior work commonly elicits Likert-scale responses or agreement levels. However, this focus on stated values overlooks a crucial dimension: the gap between what LLMs say and how they act.
Value-Action Consistency Study. In social science, this discrepancy—known as the value-action gap—is well documented Godin et al. (2005); Chung and Leung (2007); Blake (1999), especially in environmental and behavioral psychology, where cognitive, contextual, and social factors are known to hinder value-consistent actions Vermeir and Verbeke (2006). Theories of reasoned action help explain and predict such gaps in humans Ajzen (1980); Kaiser et al. (1999). Emerging research starts to evaluate consistency between self-report values and LLM actions, including ValueBench Ren et al. (2024) and GPV Ye et al. (2025b). While these works provided valuable evidence into value-action discrepancies, there lacks a context-aware evaluation framework and supporting dataset to systematically analyze these value-action inconsistencies across diverse situations. Our study investigated the value-action gap systematically across 132 scenarios, provided a context-aware dataset to support it, and introduced a set of alignment metrics to quantitatively measure this inconsistency.
@@ 3.1 Contextualizing Values into Scenarios
Together, these scenarios and values yield 14,784 contextualized Value-Informed Actions (VIA) dataset to assess the alignment (Table 1), involving both value-informed actions and associated explanations. We next introduce the detailed process and validation of the dataset generation.
@@ 3.2 Generate Value-Informed Actions with Explanations
Step3: Cross-Cultural Human Evaluation of the VIA Dataset. Using the optimal prompt selected by AI experts, we generated the “Value-Informed Actions (VIA)” dataset, comprising 14,784 value-informed actions contextualized across various scenarios (Table 1). To further evaluate dataset quality, we recruited 27 annotators with relevant cultural backgrounds through Prolific Prolific (2024). These annotators evaluated 90 randomly sampled actions and explanations using the same metrics as in Step 2. Each data instance was reviewed by three annotators, with majority voting used to finalize the assessments. The evaluation results are summarized in Table 2, with fine-grained performance for each culture in Appendix C. We proactively addressed confounding variables by rigorously validating the generated data to ensure its quality, with more details in Appendix C.
@@
North America Europe Aus Asia Africa
@@
Llama-3.3-70B 0.51 0.49 0.49 0.44 0.52 0.51 0.38 0.39 0.39 0.38 0.42 0.30
Gemma-2-9b 0.46 0.50 0.43 0.51 0.45 0.52 0.46 0.46 0.37 0.46 0.45 0.46
GPT-3.5-turbo 0.17 0.19 0.18 0.20 0.20 0.17 0.18 0.17 0.16 0.14 0.18 0.21
GPT-4o-mini 0.67 0.59 0.56 0.65 0.57 0.62 0.49 0.54 0.47 0.54 0.57 0.51
Deepseek-r1 0.59 0.51 0.52 0.52 0.51 0.56 0.41 0.46 0.52 0.42 0.58 0.49
Claude-sonnet-4 0.46 0.40 0.50 0.47 0.50 0.41 0.40 0.32 0.31 0.36 0.41 0.37
GPT-4o 0.53 0.54 0.53 0.51 0.53 0.53 0.49 0.47 0.40 0.50 0.44 0.44
Table 3: Averaged Value-Action Alignment Rates (i.e., F1 Scores) across 12 countries (top) and 11 social topics (bottom). The cell colors transition from bottom to top performances compared with other models.
@@ 4 Experimental Settings
We evaluated the value-action alignment of seven diverse models, spanning closed-source (i.e., GPT-4o-mini, GPT-4o Achiam et al. (2023) and GPT-3.5-turbo Ouyang et al. (2022)) and open-source (i.e., Gemma-2-9B Team (2024), Llama-3.3-70B Touvron et al. (2023), Deepseek-r1-distill-llama-70b DeepSeek-AI (2025)) models. We select these LLMs to represent state-of-the-art LLMs released from various countries. All models use a temperature $\tau=0.2$ following prior research Dammu et al. (2024) [^5] Robustness Test: we conducted experiments with 10 generations per prompt (temperature=0.2) on a data subset and found minimal variation (¡ 5%) in responses. For each of Task1 and Task2, we use eight distinct prompts following the approach in Figure 3. We average the eight responses to arrive at the final result. Task1 and Task2 are performed independently for each LLM in evaluating the alignment.
@@ 5.1 Value-Action Alignment Rates
Table 3 illustrates the value-action alignment rates differ by countries (See the social topic-wise alignment rates performance in Table 13). Among the six models, we observe that GPT4o-mini performed the mostly best with an F1 score of 0.564 (in summary). In comparison, GPT3.5-turbo performed significantly worse with the lowest score among all models at 0.179 (in summary). Grouping countries by geographic regions, we observe that LLMs tend to display a lower alignment rate in Africa and Asia compared to North America and Europe in GPT4o-mini, Deepseek, and Llama. Similarly, we also find the alignment rates vary across social topics, such as Leisure and Health topics (Table 13). These findings demonstrate that the alignment rates of LLMs are suboptimal, and vary dramatically by scenarios and models. We further computed the cross-task inconsistency analysis in Table 4.
Class Imbalance and Scoring Convention. Note that we follow established value evaluation practices Ren et al. (2024); Ye et al. (2025b) by coding "agree"=0 and "disagree"=1. With approximately 70-80% of responses being "agree" (0), the positive class ("disagree"=1) represents only 20-30% of the data, as demonstrated in Figure 4. In highly imbalanced binary classification, F1 scores naturally trend lower than 0.5. For our class distribution ( 30% positive class), a random classifier would achieve F1 $\simeq$ 0.3, making our observed scores of 0.2-0.6 mostly above random performance. Our scores are consistent with prior literature that also show challenges of value alignment in LLMs Cahyawijaya et al. (2024), where rigorous alignment evaluations typically yield scores below 0.5. This reflects the inherent difficulty of value alignment—a fundamental open problem in LLM safety research.
Model (A,D) (D,A) Total Misaligned
GPT-4o-mini 415 729 1,144 (15.48%)
GPT-3.5-turbo 36 1,385 1,421 (19.22%)
Llama-3.3-70B 802 284 1,086 (14.69%)
Gemma-2-9b 497 695 1,192 (16.13%)
Deepseek-r1 789 413 1,202 (16.26%)
Claude-Sonnet-4 903 203 1,106 (14.96%)
GPT-4o 626 427 1,053 (14.25%)
Table 4: Number of samples for cross-task inconsistency in both and total cases. (A,D) indicates Task1 is “Agree” and Task2 is “Disagree”. (D,A) means Task1 is “Disagree” and Task2 is “Agree”. The Total Misalignment is computed by (out of 7392).
@@ 8 Conclusion
Acknowledgments
This paper was supported by the Office of Naval Research Young Investigator Award and the NIH grant DA056725-01A1. We thank the reviewers, the area chair, and members of the SCALE lab at the University of Washington for their feedback on this work.
@@ Appendix B Prompt Variation Design
Rationales for Task 2 Design as a binary task. The binary choice design is methodologically essential in this work’s setting for several reasons:
  • Theoretical alignment with value-action gap literature: The binary paradigm directly operationalizes established psychological theory from  Godin et al. (2005) and  Chung and Leung (2007), where value-action gaps are measured through discrete alignment/misalignment rather than gradations. This enables direct comparison with foundational research.
  • Measurement validity requirements: To systematically evaluate value-action gaps, we need commensurable outputs between Task 1 (agree/disagree value statements) and Task 2 (value-aligned actions). Binary choices enable precise F1, distance, and ranking metrics that would be impossible with gradient responses.
  • Avoiding confounding complexity: Gradient scoring would require subjective interpretation of "degrees of value alignment," introducing annotator bias and cultural interpretation differences that would undermine our cross-cultural validity. Binary choices eliminate this subjectivity.
  • Framework extensibility: Our framework can accommodate gradient evaluation in future work by extending the action generation process to include multiple intensity levels per value orientation or chain-of-thought, while maintaining the core binary alignment assessment for systematic measurement.
Prompt Sensitivity Analysis. We proactively addressed prompt sensitivity through systematic design: eight prompt variants created via paraphrasing contexts, reordering options, and altering requirements. Responses are averaged across all variants to minimize bias. We further conducted the prompt sensitive experiments and report the detailed numbers in Table 8. As requested, we computed the prompt agreement (#mode-of-reponses / #all-responses) across all scenarios and all values. The results indicate a high agreement across all models, where 7 out of 8 prompts agreed with each other in most cases.
Model Prompt Agreement Rate
Task 1 Task 2
GPT-4o-mini 0.946 0.914
GPT-3.5-turbo 0.980 0.900
Llama 0.951 0.935
Gemma 0.967 0.903
Deepseek 0.911 0.925
Claude 0.968 0.893
GPT-4o 0.956 0.940
Table 8: The Prompt Agreement Rates across eight variants in Task 1 and Task2.
@@ Appendix C Human Annotation on Data Generation
Generated Dataset Validation We proactively addressed confounding variables through rigorous validation:
  • Human-in-the-loop quality control: Our three-step validation process (expert prompt selection → cross-cultural annotation → quality assessment) specifically evaluated whether generated actions accurately reflect target values. Expert evaluation achieved 94% sufficiency and annotators confirmed 89% correctness across cultures.
  • Robust prompt design: We generated actions using eight prompt variants (paraphrasing, reordering, requirement changes) and selected optimal prompts through systematic expert evaluation with substantial inter-rater reliability (Cohen’s Kappa = 0.7073).
  • Ablation evidence: The requested ablation study is effectively built into our design - our comparison of stated values (Task 1) vs. action selection (Task 2) isolates value-related variables from prompt artifacts. The systematic misalignment patterns across different cultural contexts and value types confirm these are genuine value-action gaps, not template artifacts.
@@ F.2 Risks in Value-Action Gaps
Individual Discrimination Unequal treatment or representation based on race, gender, etc.
@@
Polarization Amplifying societal divisions by aligning action with inconsistent stances.
Undermining Institutions Acting against values like justice or legality while claiming fairness.